OpenStudy (anonymous):

wendy took a trip from one city to another, a distance of 580 mi. she traveled part of the way by bus, which arrived at the train station just in time for wendy to complete her journey by train. the bus averaged 40 mi/h and the train averaged 65 mi/h the entire trip took 9 1/2h how long did wendy spend on the train

OpenStudy (dumbcow):

t1 +t2 = 9.5 40t1 +65t2 = 580 solve system using substitution ---> t2 = 9.5 - t1 40t1 +65(9.5 - t1) = 580

OpenStudy (anonymous):

once I find t1 I just subtract that from 9.5 to find my final answer correct?

OpenStudy (dumbcow):

yes :)

OpenStudy (anonymous):

alright so just to check for t1 I got 1.5 so the final answer would be t2=8

OpenStudy (anonymous):

thank you !

OpenStudy (dumbcow):

sorry ...and yes you are correct

OpenStudy (anonymous):

alright so just to check for t1 I got 1.5 so the final answer would be t2=8